silvernight269:

An angle is 12 more than 2 times its supplement. Determine the measure of each angle.

silvernight269:

@dude

silvernight269:

@Shadow

Shadow:

Y is 5 more than 4 times of x Y = 4x +5

Shadow:

Supplementary means it adds up to 180

Shadow:

Do you see how to do this one?

silvernight269:

yea

silvernight269:

so its 12+2x=180

Shadow:

no

Shadow:

Using my example from earlier. Lets say angle y and angle x are supplementary (they add up to 180). y is 5 more than 4 times of x. Therefore y = 4x + 5 y + x = 180 x + 4x + 5 = 180

Shadow:

or if it looks better this way.. y + x = 180 4x + 5 + x = 180

silvernight269:

so y=2x+12 x+y=180

Shadow:

Correct

silvernight269:

then how did ur equation equal 4x+5+X=180

Shadow:

You can't solve for a variable if there is another variable. So you define that variable. For my example, y = 4x + 5 So in x + y = 180, I can submit 4x + 5 for y

Shadow:

It's called substitution in math.

silvernight269:

ah

silvernight269:

so its actually 2x+x+12=180

Shadow:

yus

silvernight269:

okie

silvernight269:

are you sure? cus i got 3x=178

Shadow:

You subtracted 2. It is a 12.

silvernight269:

ohhh oof. my mistake. 168

silvernight269:

ok got it. x=56 and y=124

Shadow:

coolio